Someone answer this(:
What is the best TIP you would give to a student who doesn't understand how to Solve Equations?
Thankyou guys. This is another 10 pt question/:

Answer by ewatrrr(10682)


Hi
My recommendation for linear equations is to 'focus' on isolating the
variable on one side of the equation and take the steps that are allowable
to accomplish that.
For example: Doing allowable steps is for the 'sole purpose' of isolating x
2(x+5) - 4 = 5x + 6(x-1) |remove parenthsis by applying distributive principle
2x + 10 - 4 = 5x + 6x - 6 |combining like terms on both sides
2x + 6 = 11x - 6 |adding 6 to both sides of equation to combine like terms
2x + 12 = 11x |subtracting 2x from both sides to combine like terms
12 = 9x |dividing both sides of equation by 9 to finish!
12/9 = x
OR x = 1 1/3